OpenStudy (anonymous):

How do i factor x^2+4x-12?

OpenStudy (luffingsails):

Note: -6 * 2 = -12. Try using that to solve.

OpenStudy (anonymous):

That didn't help at all

OpenStudy (anonymous):

well i would start with finding the factors of 12

OpenStudy (anonymous):

would it be (4x-6)(4x+2)?

OpenStudy (anonymous):

Do the 4's cancel out?

OpenStudy (luffingsails):

The form that you are looking for is: (x + a) ( x - b) Where a * b = -12 and a + 6 = +4 So, (x + 6) (x - 2) = x^2 + 4x - 12

OpenStudy (anonymous):

after looking at the factors 1,12 2,6 3,4 find which one subtracted from the other equals the middle term 4

OpenStudy (anonymous):